# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v2

@@ 4 Experiments
To demonstrate the compositionality of net fragments, we train (updating $\boldsymbol{W}^{(L)}$ with Hebbian plasticity) the model on binary images containing straight-line patterns. For evaluation (after training and without updating any connections), we use images of kinked lines, digits, characters, and line drawings of objects such as houses and clocks. These complex patterns can be expressed as compositions of straight-line net fragments, even though the CNA was not explicitly trained on such patterns. This flexible compositional behavior contrasts strongly with associative memory (see Section 2.1). To highlight the differences between ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
Besides compositionality, we introduce noise to examine the robustness and persistence of coherent structures with statistical relevance. Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects, i.e., the efficacy of dealing with subtractive noise. We introduce additive Gaussian noise to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
Further experimental details are provided in the Appendix: We describe the training and testing datasets in more detail in Appendix E, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix F, the training details of our CNA model are in Appendix G, and the training details of the autoencoder in Appendix H.
@@ 5 Results
Figure 3: Net formed by activated neurons. The left of the image shows the neuronal activation $\boldsymbol{y}^{(S2)}$ when observing a line, and the right visualizes the net formed by these neurons. For convenience, the neurons are rearranged in a circle, and the connections of two randomly selected neurons, A and B, are visualized in colors (purple and red).
We empathize that this work focuses on learning a new neural code and, in the following, demonstrates it on a rather simple dataset.
Although our result figures display activated neurons, we would like to emphasize that these neurons are carriers of connections and form part of nets. To make this point for once, Figure 3 depicts this connectivity of activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged as a circle for convenience).
@@ 5.1 Compositionality
The model weights, $\boldsymbol{W}^{(S2)}$, are trained using straight lines (see training data at the top of Figure 4), allowing the model to capture the underlying structure of straight lines within net fragments. The model composes these fragments dynamically into nets to represent much more complex structures (see validation data in rows 1-8 of Figure 4). For instance, despite not encountering these images during training, the CNA model can represent kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows $5$-$8$).
The motivation for why we build net fragments ($\boldsymbol{y}^{(S2)}$) and not just use the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise. This noise corresponds to invalid patterns not observed during training, which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons. Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
@@ 5.2 Filtering Gaussian Noise
Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to suppression of neurons not supported by recurrent connections of net fragments. In each image, the background is shown in black, and the four different features in distinct colors. The plots are created using a bias $b^{(S2)}=0.5$ and an inhibition coefficient $\gamma=1.2+0.2t$.
The illustration in Figure 5 visually demonstrates the system’s effective capability of filtering out isolated noise or small clusters thereof. Noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix E) is introduced in an area’s afferent input, from which a significant portion is removed when constructing net fragments. While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated. As additional noise is introduced, more undesired neurons receive support and persist in their activity. However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$. In Appendix I, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
Figure 6: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed. However, when inhibition increases over time, even more noise is filtered out in the course of iteration. The plot is created after adding $10\%$ noise to each feature channel (sing a bias $b^{(S2)}=0.7$ and an inhibition coefficient $\gamma=1.2+0.2t$).
Figure 6 depicts this systematic filtering over time. The additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time. The figure illustrates a substantial noise reduction already at time step $t=0$ due to the filtering by matrix $\boldsymbol{W}^{(F)}$ and initial inhibition. In each subsequent time step, additional clutter is removed. This reduction is due to increased inhibition and falling support as a consequence of the drop-out of supporting neurons, demonstrating how overlapping fragments build a coherent net of mutually supporting neurons.
@@ 5.3 Reconstructing Figures
Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, denoted in the top left corner of the panels (left column), corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$. Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$displaying the ‘activation probabilities’ (on the right) after the first time step.
If a substantial number of neurons activate and support other neurons within a net, inactive neurons can receive significant support, leading them to switch on and encouraging figure reconstruction. Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns. This reconstruction works only reliably for up to $3$ missing pixels (see Appendix I). In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
@@
When dealing with missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected. Consequently, the network can, to some extent, handle this ambiguity and model it based on the activation map, as shown in Figure 7B. Modeling this ambiguity enables having different internal interpretations, such as whether ‘it is an interrupted line that should be reconstructed’ or ‘there are two co-linear line segments’ and selecting one of them based on the threshold $b^{(S2)}$.
Increasing Inhibition Over Time Stabilizes Patterns.
Figure 8: Reconstruction of a horizontal line interrupted by $7$ pixels: The feature activation and the initial output activity at $t=0$ are interrupted but reconstructed within the two subsequent time steps ($t=1,2$). Afterward, some artifacts triggered at line ends are reduced due to increasing inhibition. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=1.8+0.1t$.
The reconstruction process unfolds across multiple sequential steps, as illustrated in Figure 8. Initially, at time step $t=0$, the net fragments represent the interrupted line. Subsequently, during the second time step ($t=1$), one of the removed pixels is reconstructed, and shortly after, at $t=2$, the entire line is represented. Thus, reconstruction can persist even amid increasing inhibition, given that inactive neurons receive adequate support. This example underscores the necessity for multiple time steps, essential not only for noise removal but also for the reconstruction of features, as neurons can undergo flipping in each time step, fostering a dynamic interplay where the flipping of neurons encourages others to follow.
@@ 5.4 Comparison to an Autoencoder
Figure 9: Comparison of the outputs generated by the proposed CNA model and an autoencoder across different tasks. Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder). Rows $1$ to $3$ illustrate how both models encode patterns that were not present in the training dataset. Rows $4$ and $5$ compare the models’ performance in denoising tasks, while row $6$ evaluates their ability to reconstruct lines.
@@
Figure 9 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix H). While the autoencoder successfully learns to accurately represent line structures in the training data - achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and decoder reconstruction - it struggles with generalization to data not observed during training. For instance, rows $1$ to $3$ demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
@@
Rows $4$ and $5$ highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced. However, as the noise level increases, its performance decreases significantly. The CNA model, by contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations. Metrics on noise reduction, precision, and recall, presented in Appendix I, further substantiate the CNA model’s superior performance.
@@
In row $6$, the models’ capabilities in reconstructing missing line segments are compared. Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed. This is confirmed by the metrics in Appendix I: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of more noise at other locations, leading to lower overall precision and recall.
@@ 6 Discussion and Outlook
The CNA presented here establishes a novel approach to construct coherent nets as neural representations of visual input, enabling arrays of elements to be treated as unified gestalts through the composability of previously learned sub-patterns (net fragments). While the concept of such net construction has been proposed theoretically (von der Malsburg et al., 2022), this work marks the first successful implementation and simulation of the idea.
Our approach offers distinct advantages over existing neural coding frameworks, particularly in terms of robustness and the ability to represent out-of-domain data. Using nets of neurons to represent sensory signals allows filtering out noise and figure completion, while using net fragments allows representing unseen objects through composing small feature patches. In contrast to past work that relied on manually constructed filter functions (e.g., Geman and Geman, 1984), the filter functions, nets in our scheme, are learned from input statistics resp. constructed on the spot, so that they will eventually be able to cover the whole gamut of patterns encountered in natural images.
An essential next step would be the construction of a system that is trained on natural images, that contains at each position of the visual field a complement of elementary feature detector neurons able to accurately represent the pixels of the input (as in Olshausen and Field, 1996), and whose recurrent connections have sufficient lateral extent to capture the most important signal correlations present in natural images. Such systems can then be expected to develop net fragments that capture and represent all textures encountered in natural images on that length scale, thus realizing the ambition expressed in Figure 1.
A first and important such operation is abstracting over position, size, and in-plane orientation of an object’s image in the input area (primary visual cortex) by the construction of an invariant net representation of the object in a separate area (which would correspond to inferotemporal cortex (Ito et al., 1995)). As hinted at in the Introduction, this can be based on structure-sensitive, ‘homeomorphic’, mappings, as proposed in von der Malsburg (1981); Anderson and Van Essen (1987); Olshausen et al. (1995); Wiskott et al. (1997); Wolfrum et al. (2008); Memisevic and Hinton (2010) or Arathorn et al. (2013). A mapping is homeomorphic if it preserves neural feature type as well as connectivity. Object classification or recognition of individual objects can then be based on a further homeomorphic mapping to an area containing a more permanent schematic representation of the object class or a sample of the object to be recognized, as realized in Wolfrum et al. (2008).
The introduction of nets as a neural code and the inclusion of alternative neurons suggest a natural basis for the mapping process by requiring inter-areal connections to be not only specific to feature type but also to alternative neuron identity. As this identity contains information on net connectivity, this specificity can implement the homeomorphy constraint of inter-areal mappings in a natural way and overcome the limitations in existing systems (Wolfrum et al., 2008).
Once the mechanism of homeomorphic mapping has been established, it is available for arbitrary novel objects, thus forming the basis for invariant object recognition at first sight (as observed in humans in Biederman and Cooper (1991)). This capability stands in stark contrast to current ANN technology, which requires invariance to be trained individually for each object, highlighting the broader generalization potential of our approach.
The simple mechanism of net formation that we model here is a first but significant step in the direction of a data architecture grounded in nets connecting sparse sets of neurons within and between areas. Nets are singled out by the properties of cooperativity and inner consistency: all neurons of an active net experience the high-order coincidence of converging excitatory signals, realizing predictive coding at each neuron. Given high sparsity of neural firing and high sparsity of connectivity, satisfying the requirement of consistency is highly non-trivial and can only be established by a learning process in which single neurons acquire converging connections coming from sets of neurons that fire simultaneously again and again, as realized in our model.
In summary, the CNA demonstrates the feasibility and potential of a principled and compositional approach to neural coding. Its ability to autonomously construct robust representations, handle out-of-domain data, and generalize to novel stimuli underscores its promise as a foundation for more advanced systems in the future, especially when extended with multiple areas.
@@
Appendix A Weight Initialization
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are copied into the $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h^{(W2)}/2,w^{(W2)}/2)$) to $1$ if $c_{in}=\lceil c_{out}/n_{a}\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $n_{a}$ the number of alternative neurons.
The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons, which is set to $1$. When Hebbian plasticity is employed on the weight matrix $\boldsymbol{W}^{(L)}$, these connections ought to grow (as neurons tend to keep their activation state between time steps), but due to the constraint of keeping all connection weights within the range $(0,1)$ (refer to equation 8), they stay put at $1$.
Appendix B Hebbian Plasticity
After processing for $T$ time steps both matrices $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ undergo Hebbian updates (Hebb, 1949):
$\boldsymbol{W}^{(S2)}:=\min\left(\max\left(\boldsymbol{W}^{(S2)}+\alpha\cdot\boldsymbol{\rho}_{avg},0\right),1\right)$ (8)
where $\alpha$ represents the learning rate. Note that the network employs shared connections, formulated as convolutional kernels, between input patches and output neurons, making the building of fragments translation equivariant. Consequently, the same recurrent connection is applied across multiple output and input neurons. The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects. If the resulting average is positive, the corresponding connection increases by $\alpha$ as it connects more simultaneously active neurons than neurons that fire disjointly. Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced by $\alpha$.
The actual modification eq. (8) is applied once after the last iteration over $T$ timesteps to increase stability against fluctuation of the activities $\boldsymbol{y}^{(S2)}[t]$ in the course of iterations.
Appendix C Normalization
We use two normalization steps to confine the neuronal activations $\boldsymbol{a}^{\prime(S2)}$ to the interval $(0,1)$.

### v3

@@ 4 Experiments
To demonstrate the compositionality of net fragments into nets, we train the model on binary images containing straight-line patterns so that the CNA learns net fragments corresponding to straight-line sub-patterns (smaller straight-line segments). However, for evaluation (after training and without weight updates), we use much more complex images such as line drawings of objects, and demonstrate that these complex patterns can be expressed as compositions (nets) of straight-line net fragments. This flexible compositional behavior contrasts strongly with conventional associative memory that cannot model such out-of-distribution data (see Section 6.1). To highlight the differences to ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
We also examine robustness by introducing noise. Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects. Further experimental details are provided in the appendix: We describe the training and testing datasets in more detail in Appendix D, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix A, the training details of our CNA model are in Appendix E, and the training details of the autoencoder in Appendix F.
@@ 5 Results
Figure 3: The response of our model to a diagonal line input is displayed as a net of activated neurons. The left of the image shows the neuronal activation $\boldsymbol{y}^{(S2)}$ when observing the line, and the right visualizes the net formed by these neurons (activations $\boldsymbol{y}^{(S2)}$ with their connections $\boldsymbol{W}^{(L)}$). For convenience, the neurons are rearranged in a circle, and the net fragment (connections) of two randomly selected neurons, $n_{i}$ and $n_{j}$, are visualized in colors (purple and red).
Although the figures illustrate individual neuron activity, it is important to clarify that our objective is not to interpret patterns merely as isolated neuron activations, but rather as nets – cohesive, interconnected clusters of active neurons. To emphasize this perspective, Figure 3 illustrates the connectivity among activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged in a circular layout for clarity), after feeding a diagonal line into the network. We select randomly two neurons $n_{i}$ and $n_{j}$ in the left panel and visualize their net fragments $\mathcal{F}(n_{i})$ and $\mathcal{F}(n_{j})$ based on $\boldsymbol{W}^{(L)}$ on the right. This figure highlights that individual neurons (e.g., neurons $n_{i}$ and $n_{j}$) are embedded within broader nets of supportive connections. Without such network-based support, these neurons would be suppressed via attenuation mechanisms and would not appear in the left panel of Figure 3, which displays neuronal activity. Consequently, the subsequent visualizations should be understood as representing supported neurons embedded within nets, even though the underlying connectivity is not explicitly shown.
@@ 5.1 Compositionality
The model weights are trained using straight lines (see training data at the top of Figure 4). During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 1-8 of Figure 4). Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 1 and 2), digits and characters (rows 3 and 4), and line drawings (rows 5–8).
These results demonstrate that while individual net fragments $\mathcal{F}_{t}$ must be learned (in our case we use Hebbian learning to train the connections within $\mathcal{F}_{t}$), the recursive composition described in eq. 9 to form $\mathcal{N}^{*}_{t}$ allows the integration of previously uncombined fragments. Consequently, the model exhibits combinatorial generalization through the structural recomposition of learned primitives.
The motivation for why we construct net fragments rather than just using the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise. This noise corresponds to invalid patterns not observed during training (i.e., are not captured in net fragments), which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons. Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
@@ 5.2 Filtering Gaussian Noise
Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to a lack of support within nets.
To further investigate noise filtering, we introduce Gaussian noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix D) to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
The results are visualized in Figure 5: While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated. As additional noise is introduced, more undesired neurons receive support and persist in their activity. However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$. In Appendix G, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
Figure 6: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed. However, when attenuation increases over time, even more noise is filtered out. The plot is created after adding $10\%$ noise to each feature channel (using a bias $b^{(S2)}=0.7$).
The reason for the filtering is that the additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time. Figure 6 depicts this systematic filtering across timesteps. The figure illustrates a substantial noise reduction already at time step $t=0$ due to initial attenuation. In each subsequent time step, additional clutter is removed. This reduction is due to increased attenuation and falling support as a consequence of the drop-out of supporting neurons.
@@ 5.3 Reconstructing Figures
Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$. Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
If a fragmet $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$). Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns. This reconstruction is reliable for up to $3$ missing pixels (see Appendix G). In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
@@
In the presence of missing pattern elements, the membrane potential map $\boldsymbol{a}^{(S2)}$ is higher at spatial locations where line features are observed and lower where they are not observed but expected (as shown in Figure 7B). Depending on the context, it may be desirable either to reconstruct the missing line segment – effectively bridging the gap between discontinuous features – or to preserve the separation between distinct line elements. This behavior can be modulated by adjusting the hyperparameter $b^{(S2)}$ in eq. 18, which can govern the balance between feature completion and preservation of discontinuity, thereby allowing the model to selectively favor line reconstruction or separation.
@@ 5.4 Comparison to an Autoencoder
Figure 8: Comparison of the outputs generated by the proposed CNA model and an autoencoder. Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder). Rows $1$ to $3$ illustrate how both models encode patterns that were not present in the training dataset. Rows $4$ and $5$ compare the models’ performance in denoising tasks, while row $6$ evaluates their ability to reconstruct lines.
@@
Figure 8 compares performance between the proposed CNA model and an autoencoder (for model details, refer to Appendix F). While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training. For instance, rows $1$ to $3$ demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
@@
Rows $4$ and $5$ highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced. However, as the noise level increases, its performance decreases significantly. The CNA model, by contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations. Metrics on noise reduction, precision, and recall, presented in Appendix G, further substantiate the CNA model’s superior performance.
@@
In row $6$, the models’ capabilities for reconstructing missing line segments are compared. Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed. This observation is confirmed by the metrics in Appendix G: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of increased noise elsewhere, leading to lower overall precision and recall for figure reconstruction.
6 Relation to Alternative Concepts
To grasp the significance of the approach, it is important to compare it to alternative approaches.
6.1 Associative Memory
Associative memory (AM) (Amari, 1972; Hopfield, 1982) and the model presented in this work share the objective of retrieving previously encountered patterns, both utilize Hebbian synaptic plasticity for encoding information. However, fundamental differences exist in their structural and functional organization. In classical AM, neurons exhibit an all-to-all connectivity pattern, leading to the representation of structures as monolithic wholes, whereas our model constrains connectivity within learned net fragments. This restriction enables a more modular representation of complex patterns by dynamically assembling net fragments into structured nets, thereby allowing the formation of representations that extend beyond the training data distribution.
These differences arise from distinct learning mechanisms. In AM, a neural pattern is stored in a single event through the uniform strengthening of connections between co-activated neurons. Our approach, however, capitalizes on the observation that a large number of non-repeating patterns (as found in natural images) contain small recurring sub-patterns, such as texture patches. These recurring elements are captured within net fragments. The learning rule defined in eq. 20 selectively reinforces connections corresponding to the most relevant sub-patterns. Specifically, frequent sub-patterns exhibit a consistently positive average correlation $\rho$, which leads to the progressive increase of their associated synaptic weights. In contrast, rare or non-informative patterns exhibit low or even negative average correlations, causing their associated weights to decay toward zero. This Hebbian mechanism effectively filters out noise and ensures that only the most salient and statistically dominant sub-patterns are preserved in the learned representation.
A fundamental limitation of AM is the interference between stored patterns unless they are statistically independent (i.e., orthogonal), meaning that minimal neuron-wise overlap exists. Sparse coding, as proposed by Tsodyks and Feigel’man (2007) or Palm (2013), mitigates interference by activating only a small subset of neurons. Our model offers a more intrinsic solution by representing each input feature using multiple alternative neurons. These neurons possess identical receptive fields (as observed by Cossell et al. (2015)), but can exhibit different excitatory outputs. This redundancy enables overlapping net fragments to reduce neuron-wise interference by utilizing distinct alternative neurons.
6.2 Boltzmann Machine
Key aspects of our model become clearer when compared to the Boltzmann Machine (BM) introduced by Hinton and Sejnowski (1983). Both models allow recurrent connections (as also proposed in Mumford (2002)), but they also exhibit significant differences. Nevertheless, as our focus is limited to one area with net fragments ($S2$) and is not the full object detection architecture as proposed by von der Malsburg et al. (2022), we highlight only select differences here.
The BM makes a clear distinction between visible and hidden (“latent”) neurons. In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by attenuation in favor of those neurons that support each other through previously learned connections. A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference being that we do not insist on symmetric connections and therefore do not derive an energy function (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the absence of up to half of the connections due to lack of symmetry does not jeopardize the stability of a state). However, even without an energy function, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
6.3 Artificial Neural Networks (ANN)
Artificial neural networks (ANNs) rely on various architectures and learning techniques, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), and autoencoders (Hinton and Salakhutdinov, 2006). These systems process sensory input through multiple layers to generate a neuronal representation. A defining characteristic of ANN systems is their strict feedforward architecture, where connections from input to representation and from representation to output are trained via backpropagation of error (Rosenblatt, 1962, p. 292).
In contrast, the neuronal representation within the CNA consists of nets that activate as a whole, with individual neurons depending on the activation of others in the same net. Thus, while ANNs rely on a global structuring of neuronal activations, the CNA relies on a local self-organizing process.
Moreover, ANNs rely on hierarchical layers where neurons respond to features of increasing complexity. The CNA replaces this hierarchy with a single layer, where net fragments of different sizes capture features of varying complexity. These fragments interconnect to form larger nets that represent entire objects, naturally solving the binding problem, whereas ANNs do not encode explicit relationships between features, leading to ambiguities such as hallucinated feature conjunctions.
Furthermore, input representation in the CNA emerges from statistical regularities in the input, making it independent of specific tasks. In contrast, ANNs bias feature extraction to serve predefined training objectives, which limits their adaptability when retraining for different functions.
@@ 7 Discussion and Outlook
The CNA presented here establishes a novel approach for constructing coherent nets as representations of input data. This framework enables the encoding of complex patterns through the composability of previously learned sub-patterns (net fragments). Compared to existing neural coding frameworks, CNA offers distinct advantages, particularly in terms of robustness – demonstrated through its ability to filter noise and reconstruct figures – and its capacity to generalize to out-of-domain data, as evidenced by its capability to represent previously unseen objects.
For future research, we outline two key directions: First, an essential next step is the extension of CNA to natural image domains. This would allow net fragments to capture and encode texture information encountered in real-world scenarios, thus realizing the ambition expressed in Figure 1. Second, the current implementation of CNA is restricted to a single processing area with net fragments. Following the theoretical framework proposed by von der Malsburg et al. (2022), the architecture could be expanded to incorporate multiple areas, thereby enabling object recognition while achieving invariance to position, scale, and in-plane rotation. Object recognition could be realized using structure-sensitive, homeomorphic mappings between nets (von der Malsburg, 1981; Anderson and van Essen, 1987; Olshausen et al., 1995; Wiskott et al., 1997; Wolfrum et al., 2008; Memisevic and Hinton, 2010; Arathorn et al., 2013), which would facilitate the mapping of input representations onto a more stable schematic representation of an object class or an exemplar of the object to be recognized. The adoption of homeomorphic mappings, as opposed to alternative methods, preserves the core principle of nets and net fragments within and across multiple processing areas, thereby maintaining structural coherence in representation.
The proposed neural coding framework not only advances representational robustness but also aligns with principles of self-organization observed in cortical processing. Its ability to learn and flexibly recombine net fragments without explicit supervision suggests a pathway toward more adaptive and interpretable neural representations. Future extensions to multi-area architectures could bridge the gap between bottom-up feature learning and top-down predictive coding. These developments position CNA as a foundational step toward more structured and generalizable neural computation.
@@
Appendix A Feature Extraction
The stage $S1$ serves to extract initial binary features from the images. The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$. The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using the heaviside function $\Theta$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 12).
Appendix B Normalization
The membrane potential in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{M}(S2)}$ to the interval $(0,1)$.